The sum of the measures of the interior angels of a regular polygon is 20 times the square of the number of sides. What kind(s) of polygon(s) could this be?
Answer by stanbon(75887)

The sum of the measures of the interior angels of a regular polygon is 20 times the square of the number of sides. What kind(s) of polygon(s) could this be?
------------------
Equation:
(n-2)*180 = 20n^2
------
20n^2 -180n + 360 = 0
-------
n^2-9n+18 = 0
(n-3)(n-6) = 0
n = 3 or n = 6
---
The polygon has 3 sides (equilateral triangle)
or it has 6 sides (hexagon).
=============================
Cheers,
Stan H.